If a sample of 25 canned goods is selected, what is the probability that the mean of the sample will be larger than 20 mg?


1
Expert's answer
2021-06-24T17:38:07-0400

Let "X" be the mean sample: "X\sim N\left(\mu =25,\:\sigma ^2=15^2\right)" Then,

"Z=\frac{x-\mu }{\sigma }\sim \:N\left(0,1\right)"


The probability that the mean of the sample will be larger than 20mg is:

"X\sim N\left(\mu =25,\:\sigma ^2=15^2\right)"

"=1-P\left(Z\le \frac{25-20}{15}\right)\approx 1-P\left(Z\le 0.3333\right)"

"\approx 0.3694"